### v1

@@ 3 Identification Strategy & Estimation
One concern is that the distribution of LGB workers is not random, and states/counties that pass anti-discrimination laws have a higher concentration of LGB workers. However, using county fixed effects will make comparisons within a given county, so San Francisco will be compared to San Francisco. A potential concern would be if there is sorting of LGB workers to areas with anti-discrimination laws protecting them, changing the composition of the area. We show that there is little evidence to support this in the robustness check section.
To interpret $\beta_{0}$ and $\beta_{1}$ as being the causal effect of enacting employment protections for sexual orientation, it must be the case that there are no other factors related to changes in the gay/lesbian-heterosexual wage differential in states that do and do not have LGBT employment protections. To accurately estimate $\beta_{0}$ and $\beta_{1}$ while considering the staggered implementation of ENDAs, we employ the methodology from (Borusyak et al., 2021). This approach uses observations only for units and periods that are not yet treated to infer the never-treated potential outcome for each treated unit.
Our identifying assumption in this framework is that, in the absence of anti-discrimination laws, the trends in labor supply and wage disparities in counties with such laws would have mirrored those without them. This difference-in-differences framework hinges on parallel trends across all groups and periods, assuming no prior anticipation of the laws. While directly testing this parallel trend assumption is challenging, examining pre-treatment trends can provide some indication of its validity. We examine the outcomes relative to the timing of the laws by estimating an event study model, where we estimate coefficients for each period relative to the treatment time. This approach allows us to assess the impact of the laws over time and ensure a thorough understanding of their effects before and after implementation.
We plot the effect of the law on the labor supply and pay gap/premium in each year relative to the start date. The effect of the law in the year before the law is enacted is normalized to zero. We report the coefficients estimated via Borusyak et al. (2021) in the main text, and standard TWFE in the Appendix.
This paper’s main contribution is applying a staggered policy design approach, providing an understanding of the average effects of anti-discrimination laws. Our approach highlights the importance of considering local contexts when assessing the efficacy of legislative measures.
@@ 4 Primary Findings
4.1 Extensive & Intensive Margins, Wages
In Table 4, we present the regression on the extensive & intensive margins of labor supply with $\beta_{1}$ giving the effect of anti-discrimination laws on gay men’s labor supply and $\beta_{0}$ giving the effect for lesbian women. Columns (1-2) show that anti-discrimination laws have a significant effect in reducing the labor force participation and employment gap for gay men. The labor force participation rate gap is reduced by 2.1 p.p., and the employment gap is reduced by 1.5 p.p. Conversely, the labor force participation and employment premium that lesbian women have over their straight counterparts is reduced by 1.7 p.p. and 2.3 p.p., respectively. Columns (3-6) present the regression results on the intensive margin of labor supply as measured by weekly hours worked and weeks worked. We find contrasting effects of anti-discrimination laws on male and female outcomes. We find that the intensive labor supply gap for gay men as measured by weekly hours worked and weeks worked is insignificantly reduced, and the intensive labor supply premium for lesbian women is significantly reduced when measured in weeks worked. We later show that the positive impact for lesbian women on weekly hours and the negative effect on weeks worked for gay men are gone once controlled for occupational fixed effects.
In Table 5, we present the regression results on the wage gap as measured by hourly wages and annual wage earnings. These variables received a $\ln(Y_{ict}+1)$ transformation, to handle zero values. [^14] We see small differences in effect sizes in a log-only specification on the employed-only population. We run the regressions jointly by sex and present the regressions showing how the sexual orientation anti-discrimination laws impact both employed and unemployed people in columns (1) and (2) and limit the regressions to only those who are employed in columns (3) and (4). Our preferred specification is to look at both employed and unemployed people in the sample as employment is endogenous. Conditioning on employed workers would create an endogeneity problem since labor force participation and employment are significantly impacted by anti-discrimination laws. Given Table 4, the sample of employed workers could be changing as a result of the law changes, which would lead to a change in the sample composition. We present the regressions limited to employed workers to replicate past work on the wage gap/premium where it is common to only examine employed workers.
Gay men experience a notable increase in hourly wages of 6.2%, which is statistically significant at the 95% confidence level. Additionally, there is a substantial increase of 22.2% concerning their annual earnings, indicating significant progress towards income equality. On the other hand, lesbian women see an insignificant change in their annual earnings premium and an insignificant reduction in their hourly wage premium. The pay gap/premium is reduced quite significantly once the sample is limited to employed workers in columns (3) and (4), which is unsurprising. The effects also turn insignificant, which shows that they come from increased labor force participation and employment rates.
4.2 Occupation Fixed Effects
LGB workers tend to sort to different occupations and industries based on sex and sexual orientation (Black et al., 2007). They show that gay men tend to sort to more female-dominated fields and lesbian women sort to more male-dominated fields relative to their heterosexual counterparts. It is typical to not consider those factors because discrimination affects occupation and industry choice, which would make occupation and industry choice endogenous.
While controlling for occupation will bias the wage gap, it can still be instructive for seeing the effect of anti-discrimination laws on the wage and labor supply gap. We run these regressions with occupation fixed effects using four-digit occupation codes in the ACS to see how the pay gap/premium changes after anti-discrimination laws are passed. Using occupation-fixed effects will examine the effect of anti-discrimination laws within an occupation. One potential mechanism for anti-discrimination laws is that gay workers sort into more in-demand occupations, lowering their relative labor supply and wage gap, while lesbian workers do the opposite. However, we show that even when controlling for occupation there is a significant change in relative labor supply for men and women in same-sex partnerships and a significant change in relative pay for women in same-sex partnerships, indicating there is a significant effect of these laws even within an occupation. We present the extensive labor supply and pay regressions with occupation fixed effects in Table 6 and 7.
In these tables, there is a smaller pay gap for gay men and a smaller pay premium for lesbian women when controlling for occupation or industry. However, qualitatively the results are similar to the main results in Table 4 and Table 5. Gay men see a significant reduction in the labor supply gap, and lesbian women see a significant reduction in their labor supply and annual wage premium. These findings suggest that anti-discrimination laws have a significant effect even within occupations. The effect does not come from workers shifting into occupations with less discrimination, but even when conditional on occupation choice, there is a significant reduction in differences in labor supply across sexual orientations for men and women.
4.3 Impacts Over Time
We use Equation (1) with the inclusion of occupation-fixed effects and conduct event study regressions via Borusyak et al. (2021). This approach allows us to analyze how the impact estimates evolve. First, we present the event study regressions for male and female extensive labor supply outcomes in Figure 3. Panel (a) illustrates the effects on labor force participation rates, whereas Panel (b) shows the effects on employment levels. The plots reveal an initial jump in both employment and labor force participation rates among gay men, a trend that appears to wane over subsequent years. This pattern hints at an early beneficial impact that seems to moderate with time. Conversely, the trajectory for women follows a different pattern, indicating initial adverse effects on employment and labor force participation rates. The pre-trends look fairly parallel, which, while not directly confirming the parallel trends assumption for the post-period, lends it greater plausibility. These visualizations indicate a temporary increase in labor force participation and employment rates for gay men following the enactment of ENDAs, contrasting with a negative impact on the same metrics for lesbian women.
We present the event study regressions for male and female pay penalty/premium outcomes in Figure 4. The figures suggest a delayed response in wage parity gains for gay employees compared to their heterosexual counterparts, with a notable but transient increase in relative gay wages observed a year post-ENDA implementation. This delayed effect may come from implementation lags and the labor market outcomes captured in the ACS, which reflect earnings over a 23-month reference period. Conversely, there is a marked decline in relative wages for lesbian women following the enactment of the law, becoming significant one year after its introduction.
@@ 5.1 Endogenous Adoption of Anti-Discrimination Laws
One concern in this analysis is the endogenous adoption of anti-discrimination laws. Anti-discrimination laws are not randomly distributed. Locally, anti-discrimination laws are concentrated in larger cities, and state laws are to be concentrated in more liberal states that presumably are more accepting of and more favorable to LGB workers. The main concern is that there is an unobservable factor like general sentiment toward LGB workers that affects both the passage of laws as well as the labor market outcomes for LGB workers. In our main regressions, we alleviate this concern by using county-year-sex fixed effects.

### v2

@@ 3 Identification Strategy & Estimation
To interpret $\beta_{1}$ as the causal effect of enacting employment protections for sexual orientation, we assume that no other factors are driving changes in the same-sex versus different-sex wage differential in states that do and do not have LGBT employment protections. One concern is that the distribution of workers in same-sex partnerships is not random, and states/counties that pass anti-discrimination laws have a higher concentration of workers in same-sex partnerships. However, with county-year fixed effects, comparisons are made among individuals within the same county in the same year, ensuring that workers in San Francisco are compared to others in San Francisco in that same year. A potential concern would be if there is sorting of workers in same-sex partnerships to areas with anti-discrimination laws protecting them, changing the composition of the area. To account for this, we remove individuals who moved in the past year as a robustness check. [^16] While we address potential sorting by excluding recent movers, unobserved migration patterns could still affect the composition of same-sex partnerships in jurisdictions with anti-discrimination laws.
Another potential concern is that same-sex marriage legalization may confound our estimates. We address this by explicitly partialing out any state-specific effects of marriage legalization for same-sex partners. In particular, we include the term $\sum_{s=1}^{S}\lambda_{s}\bigl{[}\mathrm{SSP}_{ipct}\times\mathrm{SSM}_{st}\times\mathbf{1}_{\text{state}(c)=s}\bigr{]}$, which absorbs any differential labor-market impacts associated with the varying timing of same-sex marriage legalizations across states. This ensures that our key coefficient, $\beta_{1}$, isolates the independent effect of the local Employment Non-Discrimination Act. While these controls mitigate potential bias, we caution that, despite our efforts, overlapping policies may still exert residual confounding influences on the estimated effects.
To address potential biases from staggered policy adoption, we employ the methodology of Borusyak et al. (2021). Unlike standard TWFE models, which may misattribute post-treatment effects from early-treated units to later-treated ones, their approach uses only untreated observations to infer counterfactual outcomes for treated units. This strategy minimizes contamination and yields a more reliable estimate of $\beta_{1}$. Our identification hinges on the parallel trends assumption, that in the absence of anti-discrimination laws, labor market trends in counties with and without such laws would have evolved similarly. [^17] The difference-in-differences strategy hinges on the parallel trends assumption across groups and periods, with no anticipation of the laws prior to their implementation. We assess this assumption using event studies that estimate the dynamic effects of the policy, allowing us to observe pre-treatment trends and validate the plausibility of our identification strategy.
We report the coefficients estimated via Borusyak et al. (2021) in the main text using a fully saturated triple-difference (DDD) specification, and standard TWFE along with other difference-in-differences (DiD) estimators in Section 5. As we show in Section 5, the estimates for same-sex and different-sex individuals estimated separately are similar across estimators.
@@ 4 Primary Findings
4.1 Extensive & Intensive Margins
In Table 4, we present regression results that examine the extensive and intensive margins of labor supply for same-sex partnerships. The coefficient $\beta_{1}$ captures the effect of anti-discrimination laws on the labor supply outcomes for men in male same-sex partnerships and for women in female same-sex partnerships.
Panel A focuses on men in male same‐sex partnerships. Columns (1) and (2) show that anti-discrimination laws reduce both the labor force participation and employment gaps by approximately 3.8 and 3.7 percentage points, which translates into reductions of about 78% and 77% respectively. These coefficients are statistically significant at the 1% level. Columns (3) through (6) address the intensive margin, indicating modest increases in weekly hours and weeks worked among all men. However, when the sample is restricted to employed men (Columns (5) and (6)), these effects become statistically insignificant. This pattern suggests that the primary channel through which these laws operate is by increasing the rate of employment.
Panel B reports analogous estimates for women in female same‐sex partnerships. In contrast, the extensive margin results for women (Columns (1) and (2)) yield statistically insignificant coefficients of $-0.37$ and $-0.50$ percentage points. Women in same‐sex partnerships exhibit higher labor force participation rates than women in different‐sex partnerships. This higher baseline may partially explain the lack of significant improvement on the extensive margin, as there is less room for gains. On the intensive margin, Column (3) shows a modest increase of 0.95 weekly hours (significant at the 10% level) for all women, while Column (4) finds a small, statistically insignificant reduction in weeks worked. When focusing on employed women (Columns (5) and (6)), we observe a statistically significant increase of 0.70 weekly hours (at the 5% level) alongside a significant decrease of 0.89 weeks worked (at the 1% level). These mixed findings suggest that anti-discrimination laws may lead employed women to work longer hours per week but in fewer weeks overall, reflecting a complex adjustment on the intensive margin.
4.2 Wages
As shown in Table 4, anti-discrimination laws significantly affect labor force participation and employment outcomes, thereby potentially altering the composition of the employed sample and raising endogeneity concerns for wage analyses. Table 5 shows regression results on the wage gap, using real hourly wages and annual earnings in 2019 dollars, measured by percentile ranks and computed jointly for men and women. As Chen and Roth (2024) argue, this robust, scale-invariant metric reduces the impact of outliers and reflects an individual’s relative wage position, allowing us to evaluate how anti-discrimination laws shift wage standings for individuals in same-sex partnerships.
For men in same-sex partnerships, the increased employment rates reported in Table 4 are accompanied by significant wage gains. Specifically, anti-discrimination laws boost hourly wages by roughly 2.7 percentiles and annual earnings by 3 percentiles on average. Furthermore, these laws raise the probability of ranking above the 25th and 50th percentiles in the hourly wage distribution by 3.6 and 3.4 percentage points, respectively, although the effect at the 75th percentile is not significant. In the annual earnings distribution, anti-discrimination laws improve the likelihood of being in the upper tiers by 3.6, 5.2, and 3.7 percentage points at the 25th, 50th, and 75th percentiles, respectively. For women in same-sex partnerships, the wage distribution remains largely unaffected by anti-discrimination laws, as evidenced by statistically insignificant estimates across the board. This finding is consistent with the muted effects on employment outcomes observed in Table 4 for women.
These findings suggest that the wage gains for men in same-sex partnerships operate primarily through increased employment. New entrants to the labor force, individuals who were previously at the bottom of the wage distribution, enter at various points, thereby shifting the overall distribution upward.
4.3 Impact Dynamics
We estimate Equation (1) using event study regressions following Borusyak et al. (2021) to trace the evolution of the impact estimates. Figure 3 displays the results for extensive labor supply outcomes: Panel (a) shows labor force participation rates, while Panel (b) presents employment levels. For men in male same‐sex partnerships, both panels reveal an immediate increase in labor force participation and employment following the enactment of ENDAs, although this positive effect wanes over time. In contrast, we see no impacts on women in same-sex partnerships. The pre-treatment trends are largely parallel, supporting the plausibility of the parallel trends assumption, yet the post-treatment dynamics diverge sharply by gender. The evidence here suggests that ENDAs lead to a temporary boost in labor market entry for men in same-sex partnerships, while not affecting employment and participation rates for women in same-sex partnerships.
Figure 4 presents the event study regressions for male and female pay disparities. The results mirror the trends observed in Figure 3, with a significant wage jump occurring one year after the law’s enactment for men in same-sex partnerships. However, the difference between the one-year and four-year post‑enactment estimates is relatively modest, suggesting that the initial wage improvement for men stabilizes and remains persistent over time. We find no statistically significant effects on wage disparities for women in any post‑ENDA period. These findings indicate that while anti‑discrimination laws yield immediate and lasting wage benefits for men in same‑sex partnerships, the impact on wage outcomes for women in same-sex partnerships is insignificant.
@@ 5 Robustness Checks
To ensure the robustness of our findings on the labor market impacts of anti-discrimination laws, we conduct a series of sensitivity analyses addressing potential sources of bias. Specifically, we examine (i) the endogeneity of law adoption using state-level support for same-sex marriage as a proxy for public sentiment, (ii) the potential confounding influence of same-sex marriage legalization, (iii) the role of migration patterns, (iv) geographic heterogeneity between metropolitan and non-metropolitan areas, (v) reporting changes associated with ACS redesign, (vi) the impact of applying sample weights, and (vii) the validity of our identification strategy using placebo tests and alternative estimators.
@@ 5.1 Endogenous Adoption of Anti-Discrimination Laws
Anti-discrimination laws are not randomly distributed. Locally, anti-discrimination laws are concentrated in larger cities, and state laws are to be concentrated in more liberal states that presumably are more accepting of and more favorable to LGB workers. The main concern is that there is an unobservable factor like general sentiment toward LGB workers that affects both the passage of laws as well as the labor market outcomes for LGB workers. In our main regressions, we alleviate this concern by using county-year fixed effects.